OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the first and second equations, then the second and third equations. 3x – 5y + 2z = 7 4x + 9y – z = 2 x – 3y + 3z = 7 11x + 13y = 11 13x + 24y = 13 11x + 13y = 11 3x – 5y = 7 x – 3y = 7 13x + 24y = 13 x – 3y = 7 3x – 5y = 7

OpenStudy (anonymous):

need help ASAP :\

OpenStudy (anonymous):

eliminating from the 1st and second gives 11x+13y=11 but i cant get the second part right

OpenStudy (anonymous):

In the second equation you have a -z and in the third equation you have a 3z. To get rid of the z, you need to make the second equation's z coefficient to be the opposite of the third equation's. What could you multiply the second equation by to get a -3 in front of the z? Then add the second and third equation.

OpenStudy (anonymous):

the second equation becomes 12x+27y-3z=6 subtracted: 13x+36y=8

OpenStudy (anonymous):

am i rite?

OpenStudy (anonymous):

oh wait...lemme try that again

OpenStudy (anonymous):

subtracted: 13x+24y=-1

OpenStudy (anonymous):

eliminate the x z in the second equation , you can find out the value for z in terms of x and y now substitute this value in 1st and third equation wyou will get 1) ans corerect

OpenStudy (anonymous):

13x+24y=13...final

OpenStudy (anonymous):

yea??